# Quantum Curriculum Learning

### v3

@@ Abstract
Quantum machine learning (QML) requires significant quantum resources to address practical real-world problems. When the underlying quantum information exhibits hierarchical structures in the data, limitations persist in training complexity and generalization. Research should prioritize both the efficient design of quantum architectures and the development of learning strategies to optimize resource usage. We propose a framework called quantum curriculum learning (Q-CurL) for quantum data, where the curriculum introduces simpler tasks or data to the learning model before progressing to more challenging ones. Q-CurL exhibits robustness to noise and data limitations, which is particularly relevant for current and near-term noisy intermediate-scale quantum devices. We achieve this through a curriculum design based on quantum data density ratios and a dynamic learning schedule that prioritizes the most informative quantum data. Empirical evidence shows that Q-CurL significantly enhances training convergence and generalization for unitary learning and improves the robustness of quantum phase recognition tasks. Q-CurL is effective with broad physical learning applications in condensed matter physics and quantum chemistry.
@@
Introduction.— In the emerging field of quantum computing (QC), there is potential to use large-scale quantum computers to solve certain machine learning (ML) problems far more efficiently than classical methods. This synergy between ML and QC has given rise to quantum machine learning (QML) [1, 2], although its practical applications remain uncertain. Classical ML traditionally focuses on extracting and replicating features based on data statistics, while QML is hoped to detect correlations in classical data or generate patterns that are challenging for classical algorithms to achieve [3, 4, 5, 6, 7]. However, it remains unclear whether analyzing classical data fundamentally requires quantum effects. Furthermore, there is a question as to whether speed is the only metric by which QML algorithms should be judged [8]. This suggests a fundamental shift: it is preferable to use QML on data that is already quantum in nature [9, 10, 11, 12, 13, 14].
Figure 1: Overview of two principal methodologies in quantum curriculum learning: (a) task-based and (b) data-based approaches. In the task-based approach, a model ${\mathcal{M}}$, designated for a main task that may be challenging or constrained by data accessibility, benefits from pre-training on an auxiliary task. This auxiliary task is either relatively simpler (left panel of (a)) or has a richer dataset (right panel of (a)). In the data-based approach, we implement a dynamic learning schedule to modulate data weights, thereby emphasizing the significance of quantum data in optimizing the loss function to reduce the generalization error.
@@
The perspective of quantum resources refocuses our attention on the concept of learning. In ML, learning refers to the process through which a computer system enhances its performance on a specific task over time by acquiring and integrating knowledge or patterns from data. We can improve current QML algorithms by making this process more efficient. For example, curriculum learning [18], inspired by human learning, builds on the idea of introducing simpler concepts before progressing to complex ones, forming a strategy—a curriculum—that presents easier samples or tasks first. Although curriculum learning has been extensively applied in classical ML [19, 20, 21], its exploration in the QML field, especially regarding quantum data, is still in the early stages. Existing research has primarily examined model transfer learning in hybrid classical-quantum networks [22], where a pre-trained classical model is enhanced by adding a variational quantum circuit. However, there is still limited evidence showing that curriculum learning can effectively improve QML by scheduling tasks and samples.
@@
We explore the potential of curriculum learning using quantum data. We implement a quantum curriculum learning (Q-CurL) framework in two common scenarios. First, a main quantum task, which may be challenging due to the high-dimensional nature of the parameter space or the limitation of data availability, can be facilitated through the hierarchical parameter adjustment of auxiliary tasks. These auxiliary tasks are comparatively easier or more data-rich. However, it is necessary to establish the criteria that make an auxiliary task beneficial for a main task. Second, QML often involves noisy inputs that exhibit a hierarchical arrangement of entanglement or noisy labels, reflecting levels of importance during the optimization process. Recognizing these levels is essential for ensuring the robustness and reliability of QML methods in practical scenarios.
@@
We propose two principal approaches to address the outlined scenarios: task-based Q-CurL [Fig. 1(a)] for the first and data-based Q-CurL [Fig. 1(b)] for the second scenario. In task-based Q-CurL, the curriculum order is defined by the fidelity-based kernel density ratio between quantum datasets. This enables efficient auxiliary task selection without solving each one, reducing data demands for the main task and decreasing training epochs, even if total data requirements stay constant. In data-based Q-CurL, we employ a dynamic learning schedule that adjusts data weights to prioritize quantum data in optimization. This adaptive cost function is broadly applicable to any cost function without requiring additional quantum resources. Empirical evidence shows that task-based Q-CurL enhances training convergence and generalization when learning complex unitary dynamics. Additionally, data-based Q-CurL increases robustness, particularly in noisy-label scenarios, by preventing complete memorization of the training data. This avoids overfitting and improves generalization in the quantum phase detection task. These results suggest that Q-CurL could be broadly effective for physical learning applications.
@@
Task-based Q-CurL.— We formulate a framework for task-based Q-CurL. The target of learning is to find a function (or hypothesis) $h:{\mathcal{X}}\to{\mathcal{Y}}$ within a hypothesis set ${\mathcal{H}}$ that approximates the true function $f$ mapping ${\bm{x}}\in{\mathcal{X}}$ to ${\bm{y}}=f({\bm{x}})\in{\mathcal{Y}}$. To evaluate the correctness of $h$ given the data $({\bm{x}},{\bm{y}})$, the loss function $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is used to measure the approximation error $\ell(h({\bm{x}}),{\bm{y}})$ between the prediction $h({\bm{x}})$ and the target ${\bm{y}}$. We aim to find $h\in{\mathcal{H}}$ that minimizes the expected risk over the distribution $P({\mathcal{X}},{\mathcal{Y}})$:
@@
In practice, since the data generation distribution $P({\mathcal{X}},{\mathcal{Y}})$ is unknown, we use the observed dataset ${\mathcal{D}}={({\bm{x}}_{i},{\bm{y}}_{i})}_{i=1}^{N}\subset{\mathcal{X}}\times{\mathcal{Y}}$ to minimize the empirical risk, defined as the average loss over the training data:
$\hat{R}(h)=\frac{1}{N}\sum_{i=1}^{N}\ell(h({\bm{x}}_{i}),{\bm{y}}_{i}).$ (2)
Given a main task ${\mathcal{T}}_{M}$, the goal of task-based Q-CurL is to design a curriculum for solving auxiliary tasks to enhance performance compared to solving the main task alone. We consider ${\mathcal{T}}_{1},\ldots,{\mathcal{T}}_{M-1}$ as the set of auxiliary tasks. The training dataset for task ${\mathcal{T}}_{m}$ is ${\mathcal{D}}_{m}\subset{\mathcal{X}}^{(m)}\times{\mathcal{Y}}^{(m)}$ ($m=1,\ldots,M$), containing $N_{m}$ data pairs. We focus on supervised learning tasks with input quantum data ${\bm{x}}^{(m)}_{i}$ in the input space ${\mathcal{X}}^{(m)}$ and corresponding target quantum data ${\bm{y}}^{(m)}_{i}$ in the output space ${\mathcal{Y}}^{(m)}$ for $i=1,\ldots,N_{m}$. The training data $\left({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i}\right)$ for task ${\mathcal{T}}_{m}$ are drawn from the probability distribution $P^{(m)}({\mathcal{X}}^{(m)},{\mathcal{Y}}^{(m)})$ with the density $p^{(m)}({\mathcal{X}}^{(m)},{\mathcal{Y}}^{(m)})$. We assume that all tasks share the same data spaces ${\mathcal{X}}^{(m)}\equiv{\mathcal{X}}$ and ${\mathcal{Y}}^{(m)}\equiv{\mathcal{Y}}$, as well as the same hypothesis $h$ and loss function $\ell$ for all $m$.
@@
$R_{T_{M}}(h)$ $={\mathbb{E}}_{({\bm{x}},{\bm{y}})\sim P^{(M)}}\left[\ell(h({\bm{x}}),{\bm{y}})\right]$
@@
The curriculum weight $c_{M,m}$ can be determined using the density ratio $r({\bm{x}},{\bm{y}})=\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}$ without requiring the density estimation of $p^{(M)}({\bm{x}},{\bm{y}})$ and $p^{(m)}({\bm{x}},{\bm{y}})$. The key idea is to estimate $r({\bm{x}},{\bm{y}})$ using a linear model $\hat{r}({\bm{x}},{\bm{y}}):={\bm{\alpha}}^{\top}{\bm{\phi}}({\bm{x}},{\bm{y}})=\sum_{i=1}^{N_{M}}\alpha_{i}\phi_{i}({\bm{x}},{\bm{y}}),$ where the vector of basis functions is ${\bm{\phi}}({\bm{x}},{\bm{y}})=(\phi_{1}({\bm{x}},{\bm{y}}),\ldots,\phi_{N_{M}}({\bm{x}},{\bm{y}}))$, and the parameter vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{N_{M}})^{\top}$ is learned from data [23].
@@
The parameter vector ${\bm{\alpha}}$ is estimated via the problem of minimizing $\dfrac{1}{2}{\bm{\alpha}}^{\top}{\bm{H}}{\bm{\alpha}}-{\bm{h}}^{\top}{\bm{\alpha}}+\dfrac{\lambda}{2}{\bm{\alpha}}^{\top}{\bm{\alpha}},$ where we consider the regularization coefficient $\lambda$ for $L_{2}$-norm of ${\bm{\alpha}}$. Here, ${\bm{H}}$ is the $N_{M}\times N_{M}$ matrix with elements $H_{ll^{\prime}}=\dfrac{1}{N_{m}}\sum_{i=1}^{N_{m}}\phi_{l}({\bm{x}}_{i}^{(m)},{\bm{y}}_{i}^{(m)})\phi_{l^{\prime}}({\bm{x}}_{i}^{(m)},{\bm{y}}_{i}^{(m)})$, and ${\bm{h}}$ is the $N_{M}$-dimensional vector with elements $h_{l}=\frac{1}{N_{M}}\sum_{i=1}^{N_{M}}\phi_{l}({\bm{x}}_{i}^{(M)},{\bm{y}}_{i}^{(M)})$.
We consider each $\hat{r}({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})$ as the contribution of the data $({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})$ from the auxiliary task ${\mathcal{T}}_{m}$ to the main task ${\mathcal{T}}_{M}$. We define the curriculum weight $c_{M,m}$ as (see [23] for more details):

### v4

@@ Abstract
Quantum machine learning (QML) requires significant quantum resources to address practical real-world problems. When the underlying quantum information exhibits hierarchical structures in the data, limitations persist in training complexity and generalization. Research should prioritize both the efficient design of quantum architectures and the development of learning strategies to optimize resource usage. We propose a framework called quantum curriculum learning (Q-CurL) for quantum data, where the curriculum introduces simpler tasks or data to the learning model before progressing to more challenging ones. Q-CurL exhibits robustness to noise and data limitations, which is particularly relevant for current and near-term noisy intermediate-scale quantum devices. We achieve this through a curriculum design based on quantum data density ratios and a dynamic learning schedule that prioritizes the most informative quantum data. Empirical evidence shows that Q-CurL significantly enhances training convergence and generalization for unitary learning and improves the robustness of quantum phase recognition tasks. Q-CurL is effective with physical learning applications in physics and quantum chemistry.
@@
I Introduction
In the emerging field of quantum computing (QC), there is potential to use large-scale quantum computers to solve certain machine learning (ML) problems far more efficiently than classical methods. This synergy between ML and QC has given rise to quantum machine learning (QML) [1, 2], although its practical applications remain uncertain. Early QML research focused on quantum algorithms that theoretically enhance the efficiency of linear algebra subroutines critical to ML. A notable example is the Harrow-Hassidim-Lloyd (HHL) algorithm [3], which is designed to solve large systems of linear equations exponentially faster than classical computers. However, the HHL algorithm’s potential relies on careful preconditioning [4] to accelerate quantum computations on qubits, without considering the time required for input/output processes. These processes involve loading classical data into quantum states and extracting classical solutions from quantum states, which can be prohibitively slow, potentially negating the quantum speedup. Furthermore, if classical algorithms can efficiently utilize computational basis measurements required by a quantum algorithm, they can also exploit these measurements to accelerate linear algebra operations, rendering computation time independent of the problem’s dimensionality. This concept, known as dequantization [5], underscores a significant challenge to achieving quantum advantage.
Classical ML traditionally focuses on extracting and replicating features based on data statistics. Inspired by the success of deep learning, QML with the central techniques like quantum circuit learning [6, 7] with variational quantum algorithms [8], quantum kernel with quantum feature maps [9, 10], and quantum reservoir computing [11] with input-driven quantum dynamics, has drawn much attention in recent years. In these approaches, the central concept is to transform the task of learning from classical data into the task of identifying distinguishing features within quantum states in Hilbert space. QML is hoped to detect correlations in classical data or generate patterns that are challenging for classical algorithms to achieve [9, 10, 12, 13, 14, 15]. However, it remains unclear whether analyzing classical data fundamentally requires quantum effects. Furthermore, there is a question as to whether speed is the only metric by which QML algorithms should be judged [16]. This suggests a fundamental shift: it is preferable to use QML on data that is already quantum in nature [17, 18, 19, 20, 21, 22].
Figure 1: Overview of two principal methodologies in quantum curriculum learning: (a) task-based and (b) data-based approaches. In the task-based approach, a model ${\mathcal{M}}$, designated for a main task that may be challenging or constrained by data accessibility, benefits from pre-training on an auxiliary task. This auxiliary task is either relatively simpler (left panel of (a)) or has a richer dataset (right panel of (a)). In the task-based approach, parameter transfer is executed by initializing the model parameters of the main task with the optimal parameters derived from the auxiliary task. In the data-based approach, we implement a dynamic learning schedule to modulate data weights, thereby emphasizing the significance of quantum data in optimizing the loss function to reduce the generalization error.
@@ I Introduction
The perspective of quantum resources refocuses our attention on the concept of learning. In ML, learning refers to the process through which a computer system enhances its performance on a specific task over time by acquiring and integrating knowledge or patterns from data. We can improve current QML algorithms by making this process more efficient. This approach aligns with the intriguing perspective that describes intelligence as the efficiency of skill acquisition [26]. Curriculum learning, inspired by human educational strategies, offers a promising framework to achieve such efficiency. Proposed by Bengio et al. [27], curriculum learning involves presenting training data or tasks in a structured order, progressing from simpler to more complex examples, thereby facilitating more effective learning. This strategy mirrors human learning, where foundational concepts are mastered before tackling advanced ones, enabling a model to build robust representations incrementally.
In classical ML, various types of curriculum learning has been widely adopted to improve training efficiency and model performance across various domains [28]. Vanilla curriculum learning, introduced by Bengio et al. [27], utilizes rule-based criteria to order samples from simple to complex, thereby improving the convergence of training. Balanced curriculum learning adds diversity constraints to ensure varied samples at each stage, preventing overfitting [29, 30]. Self-paced curriculum learning, introduced by Jiang et al. [31], combines predefined and learning-based criteria for tasks like matrix factorization and multimedia event detection. Progressive curriculum learning applies curriculum concepts to model capacity or task settings, as seen in Karras et al. [32] for growing Generative Adversarial Networks. Teacher-student curriculum learning, proposed by Matiisen et al. [33], uses an auxiliary model to guide the primary model’s learning parameters, optimizing training policy. Implicit curriculum learning, exemplified by Sinha et al. [34], integrates curriculum effects indirectly, such as gradually deblurring activation maps, enhancing complexity without explicit sample ordering. Techniques such as data parameters [35], loss-based sample weighting [36] and dynamic curriculum scheduling [37] further enhance training by adaptively prioritizing samples based on their difficulty or relevance. These approaches have been successfully applied in computer vision to prioritize simpler images or features during training [38], in natural language processing to sequence samples by difficulty [39], and in reinforcement learning to guide agents through progressively challenging environments [40].
Although curriculum learning has been extensively applied in classical ML, its exploration in the QML field, especially regarding quantum data, is still in the early stages. Existing research has primarily examined model transfer learning in hybrid classical-quantum networks [41], where a pre-trained classical model is enhanced by adding a variational quantum circuit. During the revision of this manuscript, we identified a relevant application of curriculum learning in quantum circuit architecture search for determining the ground state of a given Hamiltonian [42]. This approach employs a warm-start strategy, initializing training with an easily prepared approximate state before refining the solution using reinforcement learning. However, there is still limited evidence showing that curriculum learning can effectively improve QML by scheduling tasks and samples.
@@
We explore the potential of curriculum learning using quantum data. We implement a quantum curriculum learning (Q-CurL) framework in two common scenarios. First, a main quantum task, which may be challenging due to the high-dimensional nature of the parameter space or the limitation of data availability, can be facilitated through the hierarchical parameter adjustment of auxiliary tasks. These auxiliary tasks are comparatively easier or more data-rich. Here, the parameters learned from the auxiliary task serve as an initial configuration for the parameters of the main task. This approach is particularly significant in the context of quantum data learning, where preparing perfect, noiseless quantum states is challenging. By leveraging noisy quantum data or easily prepared quantum states in auxiliary tasks, the learning process for the main task can be accelerated, even with a limited amount of training data. However, it is necessary to establish the criteria that make an auxiliary task beneficial for a main task. Second, QML often involves noisy inputs that exhibit a hierarchical arrangement of entanglement or noisy labels, reflecting levels of importance during the optimization process. Recognizing these levels is essential for ensuring the robustness and reliability of QML methods in practical scenarios.
@@
We propose two principal approaches to address the outlined scenarios: task-based Q-CurL [Fig. 1(a)] for the first and data-based Q-CurL [Fig. 1(b)] for the second scenario. In task-based Q-CurL, the curriculum order is defined by the quantum-based kernel density ratio between quantum datasets. This enables efficient auxiliary task selection without solving each one, reducing data demands for the main task and decreasing training epochs, even if total data requirements stay constant. In data-based Q-CurL, we employ a dynamic learning schedule that adjusts data weights to prioritize quantum data in optimization. This adaptive cost function is broadly applicable to any cost function without requiring additional quantum resources. Empirical evidence shows that task-based Q-CurL enhances training convergence and generalization when learning complex unitary dynamics. Additionally, data-based Q-CurL increases robustness, particularly in noisy-label scenarios, by preventing complete memorization of the training data. This avoids overfitting and improves generalization in the quantum phase detection task. These results suggest that Q-CurL could be broadly effective for physical learning applications.
II Method
II.1 Task-based Q-CurL
We formulate a framework for task-based Q-CurL. In classical ML, it is well-known that learning from multiple tasks can lead to better and more efficient algorithms. This idea encompasses areas such as transfer learning, multitask learning, and meta-learning, all of which have significantly advanced deep learning. Unlike classical ML, which typically assumes a fixed amount of training data for all tasks, in quantum learning, the order of tasks and the allocation of training data to each task are even more critical. Properly scheduling tasks could reduce the resources required for training the main task, bringing QML closer to practical applications.
@@ II.1 Task-based Q-CurL
The target of learning is to find a function (or hypothesis) $h:{\mathcal{X}}\to{\mathcal{Y}}$ within a hypothesis set ${\mathcal{H}}$ that approximates the true function $f$ mapping ${\bm{x}}\in{\mathcal{X}}$ to ${\bm{y}}=f({\bm{x}})\in{\mathcal{Y}}$. To evaluate the correctness of $h$ given the data $({\bm{x}},{\bm{y}})$, the loss function $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is used to measure the approximation error $\ell(h({\bm{x}}),{\bm{y}})$ between the prediction $h({\bm{x}})$ and the target ${\bm{y}}$. We aim to find $h\in{\mathcal{H}}$ that minimizes the expected risk over the data generation distribution $P({\mathcal{X}},{\mathcal{Y}})$:
@@
In practice, since $P({\mathcal{X}},{\mathcal{Y}})$ is unknown, we use the observed dataset ${\mathcal{D}}={({\bm{x}}_{i},{\bm{y}}_{i})}_{i=1}^{N}\subset{\mathcal{X}}\times{\mathcal{Y}}$ to minimize the empirical risk, defined as the average loss over the training data:
$\hat{R}(h)=\frac{1}{N}\sum_{i=1}^{N}\ell(h({\bm{x}}_{i}),{\bm{y}}_{i})=\frac{1}{N}\sum_{i=1}^{N}\ell_{i},$ (2)
where $\ell_{i}=\ell(h({\bm{x}}_{i}),{\bm{y}}_{i})$ is the single loss corresponding with the training data $({\bm{x}}_{i},{\bm{y}}_{i})$.
In our study, consistent with traditional approaches, the hypothesis set ${\mathcal{H}}$ is defined as as a parametric collection of hypotheses, denoted as as =${\mathcal{H}}=\{h_{{\bm{\theta}}}\}$, where ${\bm{\theta}}$ represents the model parameters, and each value of ${\bm{\theta}}$ corresponds to a specific hypothesis $h_{{\bm{\theta}}}$. The objective is to initialize the model parameters at an appropriate starting point, ${\bm{\theta}}={\bm{\theta}}_{0}$, and iteratively optimize them to identify the optimal parameters, ${\bm{\theta}}={\bm{\theta}}_{\textup{opt}}$, that minimize the empirical risk $\hat{R}(h_{\bm{\theta}})$.
II.1.1 Design a curriculum via curriculum weights
Given a main task ${\mathcal{T}}_{M}$, the goal of task-based Q-CurL is to design a curriculum for solving auxiliary tasks to enhance performance compared to solving the main task alone. We consider ${\mathcal{T}}_{1},\ldots,{\mathcal{T}}_{M-1}$ as the set of auxiliary tasks. The training dataset for ${\mathcal{T}}_{m}$ is ${\mathcal{D}}_{m}\subset{\mathcal{X}}^{(m)}\times{\mathcal{Y}}^{(m)}$ ($m=1,\ldots,M$), containing $N_{m}$ data pairs. We focus on supervised learning tasks with input quantum data ${\bm{x}}^{(m)}_{i}$ in the input space ${\mathcal{X}}^{(m)}$ and corresponding target data ${\bm{y}}^{(m)}_{i}$ in the output space ${\mathcal{Y}}^{(m)}$ for $i=1,\ldots,N_{m}$. The training data $\left({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i}\right)$ for ${\mathcal{T}}_{m}$ are drawn from the probability distribution $P^{(m)}({\mathcal{X}}^{(m)},{\mathcal{Y}}^{(m)})$ with the density $p^{(m)}({\mathcal{X}}^{(m)},{\mathcal{Y}}^{(m)})$. We assume that all tasks share the same data spaces ${\mathcal{X}}^{(m)}\equiv{\mathcal{X}}$ and ${\mathcal{Y}}^{(m)}\equiv{\mathcal{Y}}$, as well as the same hypothesis class $\{h_{{\bm{\theta}}}\}$ and the same loss function $\ell$ for all $m$.
We focus on identifying an auxiliary task ${\mathcal{T}}_{m}$ such that solving ${\mathcal{T}}_{m}$ facilitates the solution of the main task ${\mathcal{T}}_{M}$. For clarity, we denote the parameters of task ${\mathcal{T}}_{m}$ as ${\bm{\theta}}^{(m)}$. In this curriculum scheme, only model parameters are transferred. Specifically, upon solving ${\mathcal{T}}_{m}$, we obtain the optimal parameter ${\bm{\theta}}^{(m)}={\bm{\theta}}^{(m)}_{\textup{opt}}$. These parameters are then used as the initial values for the main task, setting ${\bm{\theta}}^{(M)}={\bm{\theta}}^{(M)}_{0}={\bm{\theta}}^{(m)}_{\textup{opt}}$, before iteratively optimizing to achieve the optimal parameters ${\bm{\theta}}^{(M)}={\bm{\theta}}^{(M)}_{\textup{opt}}$ for ${\mathcal{T}}_{M}$.
@@ II.1.1 Design a curriculum via curriculum weights
$R_{T_{M}}(h)$ $={\mathbb{E}}_{({\bm{x}},{\bm{y}})\sim P^{(M)}}\left[\ell(h({\bm{x}}),{\bm{y}})\right]$ (3)
$=\int\int_{({\bm{x}},{\bm{y}})}\ell(h({\bm{x}}),{\bm{y}})p^{(M)}({\bm{x}},{\bm{y}})d({\bm{x}},{\bm{y}})$ (4)
$=\int\int_{({\bm{x}},{\bm{y}})}\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}\ell(h({\bm{x}}),{\bm{y}})p^{(m)}({\bm{x}},{\bm{y}})d({\bm{x}},{\bm{y}})$ (5)
@@
The curriculum weight $c_{M,m}$ can be determined using the density ratio $r({\bm{x}},{\bm{y}})=\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}$ without requiring the density estimation of $p^{(M)}({\bm{x}},{\bm{y}})$ and $p^{(m)}({\bm{x}},{\bm{y}})$. Similar to the unconstrained least-squares importance fitting approach [43] in classical ML, the key idea is to model the density ratio function $r({\bm{x}},{\bm{y}})$ using a linear model:
$\hat{r}_{{\bm{\alpha}}}({\bm{x}},{\bm{y}}):={\bm{\alpha}}^{\top}{\bm{\phi}}({\bm{x}},{\bm{y}})=\sum_{i=1}^{N_{M}}\alpha_{i}\phi_{i}({\bm{x}},{\bm{y}}),$ (7)
where the vector of basis functions is ${\bm{\phi}}({\bm{x}},{\bm{y}})=(\phi_{1}({\bm{x}},{\bm{y}}),\ldots,\phi_{N_{M}}({\bm{x}},{\bm{y}}))$, and the parameter vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{N_{M}})^{\top}$ is learned from data.
The basis function $\phi_{l}({\bm{x}},{\bm{y}})$ is defined as the product of kernels used to compare two pairs of input and output states as:
$\phi_{l}({\bm{x}},{\bm{y}})={\mathcal{K}}_{x}[{\bm{x}}{\bm{x}}^{(M)}_{l}]{\mathcal{K}}_{y}[{\bm{y}}{\bm{y}}^{(M)}_{l}].$ (8)
Here, ${\mathcal{K}}_{x}(\cdot,\cdot)$ and ${\mathcal{K}}_{y}(\cdot,\cdot)$ are the kernels defined in the data space ${\mathcal{X}}$ and ${\mathcal{Y}}$. The key factor that differentiates this framework from classical curriculum learning is the consideration of quantum data for ${\bm{x}}$ and ${\bm{y}}$, which are assumed to be in the form of density matrices representing quantum states. For example, the kernel function ${\mathcal{K}}_{x}$ and ${\mathcal{K}}_{y}$ can be naturally defined as the global fidelity kernel, which leads to the form of $\phi_{l}({\bm{x}},{\bm{y}})$ as
@@
In our numerical experiments, which focus on learning unitary dynamics using quantum data for both input and output, fidelity emerges as an appropriate metric due to its direct relevance to the task. However, to enhance scalability for large-scale quantum data, efficient quantum kernels such as the quantum projected kernel [44] or the shadow tomography kernel [45] can be utilized to estimate the density ratio effectively. Moreover, in other ML scenarios such as classifying quantum data with quantum inputs and classical outputs, a hybrid approach may prove more effective. In these cases, a quantum kernel can be applied to the quantum components, while a classical kernel complements the classical parts.
In this way, $R_{T_{M}}(h)$ can be approximated by
$R_{T_{M}}(h)\approx{\mathbb{E}}_{({\bm{x}},{\bm{y}})\sim P^{(m)}}\left[\hat{r}_{\bm{\alpha}}({\bm{x}},{\bm{y}})\ell(h({\bm{x}}),{\bm{y}})\right],$ (10)
or, as an approximation, using the following sample averages:
@@
The parameter vector ${\bm{\alpha}}$ is estimated by minimizing the following error:
[…]
Given the training data, we can further reduce the minimization of Eq. (13) to the problem of minimizing
$\dfrac{1}{2N_{m}}\sum_{i=1}^{N_{m}}\hat{r}_{\bm{\alpha}}^{2}({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})-\dfrac{1}{N_{M}}\sum_{i=1}^{N_{M}}\hat{r}_{\bm{\alpha}}({\bm{x}}^{(M)}_{i},{\bm{y}}^{(M)}_{i})+\dfrac{\lambda}{2}\|{\bm{\alpha}}\|_{2}^{2},$ (14)
where we consider the regularization coefficient $\lambda$ for $L_{2}$-norm of ${\bm{\alpha}}$. Equation (14) can be further reduced to the following quadratic form:
$\min_{{\bm{\alpha}}}\dfrac{1}{2}{\bm{\alpha}}^{\top}{\bm{H}}{\bm{\alpha}}-{\bm{h}}^{\top}{\bm{\alpha}}+\dfrac{\lambda}{2}{\bm{\alpha}}^{\top}{\bm{\alpha}}.$ (15)
Here, ${\bm{H}}$ is the $N_{M}\times N_{M}$ matrix with elements $H_{ll^{\prime}}=\dfrac{1}{N_{m}}\sum_{i=1}^{N_{m}}\phi_{l}({\bm{x}}_{i}^{(m)},{\bm{y}}_{i}^{(m)})\phi_{l^{\prime}}({\bm{x}}_{i}^{(m)},{\bm{y}}_{i}^{(m)})$, and ${\bm{h}}$ is the $N_{M}$-dimensional vector with elements $h_{l}=\frac{1}{N_{M}}\sum_{i=1}^{N_{M}}\phi_{l}({\bm{x}}_{i}^{(M)},{\bm{y}}_{i}^{(M)})$.
We can consider each $\hat{r}({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})$ in Eq. (11) as the contribution of the data $({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})$ from the auxiliary task ${\mathcal{T}}_{m}$ to the main task ${\mathcal{T}}_{M}$. From Eq. (11), we note that only the quantity $\ell(h({\bm{x}}^{(m)}_{i}),{\bm{y}}^{(m)}_{i})$ depends on the training performance of the auxiliary task ${\mathcal{T}}_{m}$. We assume that the loss $\ell(h({\bm{x}}^{(m)}_{i}),{\bm{y}}^{(m)}_{i})$ is bounded by a quantity $\ell_{\textup{max}}^{(m)}$ for all $i=1,\ldots,N_{m}$. Then the empirical risk $R_{T_{M}}(h)$ (before solving ${\mathcal{T}}_{M}$) can be bounded by the following inequality:
$R_{T_{M}}(h)$ $\approx\dfrac{1}{N_{m}}\sum_{i=1}^{N_{m}}\hat{r}_{\bm{\alpha}}({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})\ell(h({\bm{x}}^{(m)}_{i}),{\bm{y}}^{(m)}_{i})$ (16)
$\leq\dfrac{\ell_{\textup{max}}^{(m)}}{N_{m}}\sum_{i=1}^{N_{m}}\hat{r}_{\bm{\alpha}}({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})=\ell^{(m)}_{\textup{max}}c_{M,m},$ (17)
where the curriculum weight $c_{M,m}$ is defined as
@@
We clarify that $c_{M,m}$ quantifies the effect of minimizing $\ell^{(m)}_{\textup{max}}$, associated with the auxiliary task ${\mathcal{T}}_{m}$, on the empirical risk in training the main task ${\mathcal{T}}_{M}$. In Eq. (17), the empirical risk is bounded with the product of $l^{(m)}_{\textup{max}}$ and $c_{M,m}$. This upper bound is independent of the optimization process for the main task ${\mathcal{T}}_{M}$. Here, $\ell^{(m)}_{\textup{max}}$ is determined by the optimization process of the auxiliary task ${\mathcal{T}}_{m}$, while $c_{M,m}$ reflects the similarity between the data domains of the auxiliary and main tasks. We compare the upper bound in two scenarios: one without solving ${\mathcal{T}}_{m}$ in advance, and another where ${\mathcal{T}}_{m}$ is solved first, with its parameters transferred to ${\mathcal{T}}_{M}$. We focus on assessing the extent to which solving $\mathcal{T}_{m}$ reduces this upper bound better. More concretely, we define $\ell^{(m)}_{1}$ and $\ell^{(m)}_{2}$ as the values of $\ell^{(m)}_{\textup{max}}$ at ${\bm{\theta}}^{(m)}={\bm{\theta}}^{(m)}_{0}$ (before solving ${\mathcal{T}}_{m}$) and at ${\bm{\theta}}^{(m)}={\bm{\theta}}^{(m)}_{\textup{opt}}$ (after solving ${\mathcal{T}}_{m}$), respectively. As $\ell^{(m)}_{\textup{max}}$ depends solely on the auxiliary task ${\mathcal{T}}_{m}$, we assume that solving each ${\mathcal{T}}_{m}$ reduces $\ell^{(m)}_{\textup{max}}$ by a consistent amount, $\ell^{(m)}_{1}-\ell^{(m)}_{2}=\Delta\ell$, for all $m$. Consequently, the reduction in the upper bound in Eq. (17) after solving ${\mathcal{T}}_{m}$ is given by $(\Delta\ell)c_{M,m}$. Thus, a large (small) $c_{M,m}$ indicates that solving ${\mathcal{T}}_{m}$ has a greater (lesser) reduction in the upper bound, resulting in a more (less) significant contribution to minimizing the empirical risk $R_{T_{M}}(h)$.
II.1.2 Unitary learning task and Q-CurL game